OpenStudy (anonymous):

(a) Suppose that a NASCAR race car is moving to the right with a constant velocity of +93 m/s. What is the average acceleration of the car? (b) Twelve seconds later, the car is halfway around the track and traveling in the opposite direction with the same speed. What is the average acceleration of the car?

OpenStudy (anonymous):

I need an idea on how to get the solution, thank you.

OpenStudy (anonymous):

a) avg acceleration is 0 since there is no change in velocity b) avg acceleration = change in velocity/time = {93-(-93)}/12 = 186/12 = 15.5 m/(s*s)

OpenStudy (anonymous):

hope that helps